OpenStudy (anonymous):

help i will give a medal starting from 105 feet away a person on a bicycle rides toward a checkpoint and then passes it.the rider is traveling at a constant rate of 35 feet per second.the distance between the bicycle and the checkpoint is given by the equation d=|105-35t| at what time is the bike 60 feet away from the checkpoint???

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (perl):

so you want to solve 60 = | 105 - 35t|

OpenStudy (anonymous):

no at what time is the bike 60 feet away from the checkpoint???

OpenStudy (anonymous):

a. 1.3 sec and 2.6 sec b. 1.3 sec and 4.7 sec c. 1.1 sec and 2.6 sec d. 4.7 sec and 9.4 sec

OpenStudy (anonymous):

please help quick. i will give a medal!

OpenStudy (anonymous):

@perl

OpenStudy (perl):

i can't give answer, but I can show you how to get the answer

OpenStudy (anonymous):

please do and quick

OpenStudy (anonymous):

hello :(

OpenStudy (anonymous):

@perl

OpenStudy (perl):

so we want to solve 60 = | 105 - 35t| remove the absolute bars , 60 = +/- (105 - 35t)

OpenStudy (perl):

positive case: 60 = (105 - 35t) add 35t to both sides 35t = 105 - 60 35t = 45 t = 45/35 = 1.2857 ~ 1.3

OpenStudy (anonymous):

omg............................

OpenStudy (anonymous):

rot in hell wingspan